Stephen Curry signed a deal with Under Armour, where UA will pay Curry $4M every year, starting from 1/1/2015, for 5 years. UA expects to make $2M, $5M, $10M, $20M and $30M profit from Curry's brand for the 5 years. Suppose the profits are earned on the first day of each year from 2016 to 2020. The discount rate is 10% per year. What was the NPV of this contract at the time of 1/1/2015?
Answer #1

The cash flows will be as follows

Year Payment for brand Profits Cash flows
1/1/2015 0 -4 -4
1/1/2016 1 -4 2 -2
1/1/2017 2 -4 5 1
1/1/2018 3 -4 10 6
1/1/2019 4 -4 20 16
1/1/2020 5 30 30

NPV as on 1/1/2015 = CF0+ sum of discounted values of cash flows

= -4 -2/1.1^1 + 1/1.1^2+ 6/1.1^3+ 16/1.1^4 + 30/1.1^5

=$29.07 million